# Bo Du1† Xuekang Zhu1,2† Xiaochen Ma3† Chenfan Qu4, 2† Kaiwen Feng1† Zhe Yang1 Chi-Man Pun5 Jian Liu2‡ Jizhe Zhou1‡

### v2

@@ 6.1 Effectiveness of Low-Level Feature Extractors
Since each domain has proposed specific feature extractors, to explore their effectiveness under a unified domain, we conduct experiments using 6 backbones combined with 4 different extractors in the shallow layer under the aforementioned IFF-Protocol setting. The extractors are BayarConv Bayar_2018 for noise aritfacts, Sobel MVSS_2021 for edge arifacts, DCT ahmed2006discrete and FPH qu2023doctamper for frequency artifacts. Details of each extractor can be found in Appendix G.1.
Results in Table 8 show the AUC differences between versions of each backbone using the four different feature extractors and those without, averaged across all test datasets for each task. All backbones except EfficientNet tan2019efficientnet show performance drops after using feature extractors, indicating that under the IFF-Protocol, where training data includes sufficient manipulation types and image quantity, models do not rely on the additional information provided by feature extractors. However, due to its lightweight nature, EfficientNet still benefits from the use of feature extractors. The results suggest that feature extractors may only be beneficial for detection on small-scale datasets, with limited manipulation types, or when using lightweight models. Details of each domain test datasets AUC scores can be found in Appendix G.2.
@@
Task Extractor = Sobel MVSS_2021 Extractor = BayarBayar_2018
@@
Extractor = FPH qu2023doctamper Extractor = DCT ahmed2006discrete
@@ 6.2.1 Cross‑Evaluation Between IMDL and Document Benchmarks
Table 4 reports the within‑domain results obtained under the original Document benchmark split, whereas Table 5 presents the cross‑domain scores produced by our newly introduced generalization protocol. Across both settings, IMDL detectors demonstrate strong competitiveness in the document forensics scenario. In the conventional split, the Cat-Net CAT-Net2022 ; qu2023doctamper ; chen2024ffdn family achieves the best average F1, confirming the merit of its hierarchical “cat‑net” paradigm. Under the more challenging cross‑domain evaluation, PSCC-Net liu2022pscc displays markedly better generalization, suggesting that progressive spatial modeling captures cues for document manipulation localization. We expect future work to further investigate the underlying mechanisms behind PSCC-Net.
@@
Following the MVSS training protocol ma2024imdl , all Document-oriented models are trained on the CASIAv2 dataset CASIA_2013 and evaluated on five standard IMDL test sets. As shown in Table 9, the dual-branch architecture of CAFTB song2025caftb achieves the best overall performance among Document models when transferred to IMDL tasks—an outcome that aligns with the design philosophy of the current SoTA model Mesorch zhu2025mesoscopic , which also emphasizes dual-branch learning.
@@
CAFTB song2025caftb (TOMM24) 0.6234 0.2557 0.6731 0.3526 0.3770 0.4564
DTD qu2023doctamper (CVPR23) 0.3535 0.1482 0.6470 0.1749 0.1811 0.3009
FFDN chen2024ffdn (ECCV24) 0.5012 0.2670 0.6085 0.2516 0.2957 0.3848
TIFDM dong2024tifdm (TCE24) 0.3675 0.2087 0.3572 0.1634 0.2333 0.2660
@@ 6.2.2 Extending IMDL Detectors to AIGC and Deepfake Benchmarks
We train each IMDL detector on the FF++‑c23 training split and evaluate on all remaining deep‑fake test sets; the scores are given in Table 10. When compared to all baselines in the deepfakebench yan2023deepfakebench , Cat-Net attains the best performance in the within‑domain setting, while Mesorch achieves the highest average accuracy in the cross‑domain evaluation, establishing new state‑of‑the‑art results in both regimes.
@@
CAT-Net CAT-Net2022 (IJCV22) 0.9855 0.8351 0.9946 0.9891 0.9905 0.9689 0.9606 0.7489 0.7511 0.8086 0.7196 0.7043 0.6626 0.9282 0.7605
IML-ViT ma2023iml (Arxiv) 0.9614 0.8483 0.9810 0.9728 0.9793 0.9206 0.9439 0.7250 0.7419 0.7770 0.7516 0.7237 0.6149 0.8718 0.7437
Mesorch zhu2025mesoscopic (AAAI25) 0.9815 0.8445 0.9889 0.9890 0.9859 0.9636 0.9589 0.7900 0.7649 0.8468 0.7806 0.7345 0.6638 0.9623 0.7918
MVSS-Net MVSS_2021 (ICCV21) 0.9723 0.8140 0.9779 0.9865 0.9868 0.9458 0.9472 0.6969 0.6839 0.7749 0.6764 0.6785 0.5841 0.8871 0.7117
PSCC-Net liu2022pscc (TCSVT22) 0.7147 0.6536 0.7751 0.6414 0.6460 0.6065 0.6729 0.4602 0.5898 0.5565 0.5945 0.5845 0.6574 0.8216 0.6092
Trufor trufor2023 (CVPR23) 0.7303 0.6848 0.8087 0.7687 0.7084 0.6326 0.7223 0.6551 0.6115 0.5908 0.6017 0.5749 0.5977 0.9526 0.6549
@@ Appendix I Broader Impacts Discussion
NeurIPS Paper Checklist
  1. 1.
    Claims
  3. Answer: [Yes]
  5. 2.
    Limitations
  7. Answer: [Yes]
  8. Justification: We discuss the limitations in the Appendix.
  9. 3.
    Theory assumptions and proofs
  11. Answer: [N/A]
  13. 4.
    Experimental result reproducibility
  15. Answer: [Yes]
  17. 5.
    Open access to data and code
  19. Answer: [Yes]
  21. 6.
    Experimental setting/details
  23. Answer: [Yes]
  25. 7.
    Experiment statistical significance
  27. Answer: [No]
  29. 8.
    Experiments compute resources
  31. Answer: [Yes]
  33. 9.
    Code of ethics
  35. Answer: [Yes]
  37. 10.
    Broader impacts
  39. Answer: [Yes]
  41. 11.
    Safeguards
  43. Answer: [N/A]
  44. Justification: This paper posed no such risks.
  45. 12.
    Licenses for existing assets
  47. Answer: [Yes]
  49. 13.
    New assets
  51. Answer: [Yes]
  53. 14.
    Crowdsourcing and research with human subjects
  55. Answer: [N/A]
  57. 15.
  59. Answer: [N/A]
  61. 16.
    Declaration of LLM usage
  63. Answer: [N/A]

### v3

@@ 6.1 Effectiveness of Low-Level Feature Extractors
Since each domain has proposed specific feature extractors, to explore their effectiveness under a unified domain, we conduct experiments using 6 backbones combined with 4 different extractors in the shallow layer under the aforementioned IFF-Protocol setting. The extractors are BayarConv Bayar and Stamm (2018) for noise aritfacts, Sobel Chen et al. (2021) for edge arifacts, DCT Ahmed et al. (2006) and FPH Qu et al. (2023) for frequency artifacts. Details of each extractor can be found in Appendix G.1.
Results in Table 8 show the AUC differences between versions of each backbone using the four different feature extractors and those without, averaged across all test datasets for each task. All backbones except EfficientNet Tan and Le (2019) show performance drops after using feature extractors, indicating that under the IFF-Protocol, where training data includes sufficient manipulation types and image quantity, models do not rely on the additional information provided by feature extractors. However, due to its lightweight nature, EfficientNet still benefits from the use of feature extractors. The results suggest that feature extractors may only be beneficial for detection on small-scale datasets, with limited manipulation types, or when using lightweight models. Details of each domain test datasets AUC scores can be found in Appendix G.2.
@@
Task Extractor = Sobel Chen et al. (2021) Extractor = BayarBayar and Stamm (2018)
@@
Extractor = FPH Qu et al. (2023) Extractor = DCT Ahmed et al. (2006)
@@ 6.2.1 Cross‑Evaluation Between IMDL and Document Benchmarks
Table 4 reports the within‑domain results obtained under the original Document benchmark split, whereas Table 5 presents the cross‑domain scores produced by our newly introduced generalization protocol. Across both settings, IMDL detectors demonstrate strong competitiveness in the document forensics scenario. In the conventional split, the Cat-Net Kwon et al. (2022); Qu et al. (2023); Chen et al. (2024) family achieves the best average F1, confirming the merit of its hierarchical “cat‑net” paradigm. Under the more challenging cross‑domain evaluation, PSCC-Net Liu et al. (2022a) displays markedly better generalization, suggesting that progressive spatial modeling captures cues for document manipulation localization. We expect future work to further investigate the underlying mechanisms behind PSCC-Net.
@@
Following the MVSS training protocol Ma et al. (2024), all Document-oriented models are trained on the CASIAv2 dataset Dong et al. (2013) and evaluated on five standard IMDL test sets. As shown in Table 9, the dual-branch architecture of CAFTB Song et al. (2025) achieves the best overall performance among Document models when transferred to IMDL tasks—an outcome that aligns with the design philosophy of the current SoTA model Mesorch Zhu et al. (2025), which also emphasizes dual-branch learning.
@@
CAFTB Song et al. (2025)(TOMM24) 0.6234 0.2557 0.6731 0.3526 0.3770 0.4564
DTD Qu et al. (2023)(CVPR23) 0.3535 0.1482 0.6470 0.1749 0.1811 0.3009
FFDN Chen et al. (2024)(ECCV24) 0.5012 0.2670 0.6085 0.2516 0.2957 0.3848
TIFDM Dong et al. (2024)(TCE24) 0.3675 0.2087 0.3572 0.1634 0.2333 0.2660
@@ 6.2.2 Extending IMDL Detectors to AIGC and Deepfake Benchmarks
We train each IMDL detector on the FF++‑c23 training split and evaluate on all remaining deep‑fake test sets; the scores are given in Table 10. When compared to all baselines in the deepfakebench Yan et al. (2023b), Cat-Net attains the best performance in the within‑domain setting, while Mesorch achieves the highest average accuracy in the cross‑domain evaluation, establishing new state‑of‑the‑art results in both regimes.
@@
CAT-Net Kwon et al. (2022)(IJCV22) 0.9855 0.8351 0.9946 0.9891 0.9905 0.9689 0.9606 0.7489 0.7511 0.8086 0.7196 0.7043 0.6626 0.9282 0.7605
IML-ViT Ma et al. (2023b)(Arxiv) 0.9614 0.8483 0.9810 0.9728 0.9793 0.9206 0.9439 0.7250 0.7419 0.7770 0.7516 0.7237 0.6149 0.8718 0.7437
Mesorch Zhu et al. (2025)(AAAI25) 0.9815 0.8445 0.9889 0.9890 0.9859 0.9636 0.9589 0.7900 0.7649 0.8468 0.7806 0.7345 0.6638 0.9623 0.7918
MVSS-Net Chen et al. (2021)(ICCV21) 0.9723 0.8140 0.9779 0.9865 0.9868 0.9458 0.9472 0.6969 0.6839 0.7749 0.6764 0.6785 0.5841 0.8871 0.7117
PSCC-Net Liu et al. (2022a)(TCSVT22) 0.7147 0.6536 0.7751 0.6414 0.6460 0.6065 0.6729 0.4602 0.5898 0.5565 0.5945 0.5845 0.6574 0.8216 0.6092
Trufor Guillaro et al. (2023)(CVPR23) 0.7303 0.6848 0.8087 0.7687 0.7084 0.6326 0.7223 0.6551 0.6115 0.5908 0.6017 0.5749 0.5977 0.9526 0.6549
@@ 8 Acknowledgment
NeurIPS Paper Checklist
  1. 1.
    Claims
  3. Answer: [Yes]
  5. 2.
    Limitations
  7. Answer: [Yes]
  8. Justification: We discuss the limitations in the Appendix.
  9. 3.
    Theory assumptions and proofs
  11. Answer: [N/A]
  13. 4.
    Experimental result reproducibility
  15. Answer: [Yes]
  17. 5.
    Open access to data and code
  19. Answer: [Yes]
  21. 6.
    Experimental setting/details
  23. Answer: [Yes]
  25. 7.
    Experiment statistical significance
  27. Answer: [No]
  29. 8.
    Experiments compute resources
  31. Answer: [Yes]
  33. 9.
    Code of ethics
  35. Answer: [Yes]
  37. 10.
    Broader impacts
  39. Answer: [Yes]
  41. 11.
    Safeguards
  43. Answer: [N/A]
  44. Justification: This paper posed no such risks.
  45. 12.
    Licenses for existing assets
  47. Answer: [Yes]
  49. 13.
    New assets
  51. Answer: [Yes]
  53. 14.
    Crowdsourcing and research with human subjects
  55. Answer: [N/A]
  57. 15.
  59. Answer: [N/A]
  61. 16.
    Declaration of LLM usage
  63. Answer: [N/A]